# BeetleVerse: A study on taxonomic classification of ground beetles

### v2

@@ 2 Related Work
The computational identification of insects has emerged as a critical research domain, with numerous studies addressing the challenges of taxonomic classification through advanced machine learning (ML) techniques [39, 57, 3, 59, 55, 5, 28]. Domain adaptation is critical [61, 49, 29, 17, 37]. For example, [49] investigated the generalization of ML models across different imaging conditions, revealing significant performance drops for out of distribution data. Several works proposed multimodal approaches integrating visual features with ecological metadata [12, 24, 2], showing different signal modalities often evolve independently to serve distinct functions. A few recent works have explored automated methods for identifying and monitoring carabids [6, 25, 21, 58, 1, 45]. Most notable of them is NEON’s continental-scale ground beetle monitoring program [25, 40]. Among others, [6] achieved 84.6% species-level accuracy using linear discriminant analysis on pitfall trap-collected carabids; [21] achieved 51.9% species-level accuracy using applied neural networks; and [1] got 90% accuracy on Sri Lankan beetles by using background removal, segmentation, and transfer learning. However, these approaches are limited by small taxonomic coverage and the use of conventional ML models that lack the representational power of cutting-edge vision encoders. Our work addresses these limitations by leveraging that power and covering a broader taxonomic range.
@@ Acknowledgment
This work is supported by both the Imageomics Institute and the AI and Biodiversity Change Global Climate Center. The Imageomics Institute is funded by the US National Science Foundation’s Harnessing the Data Revolution program under Award #2118240 (Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning). The ABC Global Climate Center is funded by the US National Science Foundation under Award No. #2330423 and Natural Sciences and Engineering Research Council of Canada under Award No. #585136. Additionally, this material is based upon work supported by the National Science Foundation under Award Numbers #2301322, #1950364, #242918, and #1926569, as well as the USDA National Institute of Food and Agriculture, Hatch Project Number ME0-22425 through the Maine Agricultural and Forest Experiment Station for SR and AE.
@@
All four datasets are publicly available: BeetlePUUM [46], BeetlePalooza [16], NHM-Carabids [22], I1MC [41].
@@
All codes to reproduce our results are available here.
@@ A.1 Dataset Details
This dataset digitize all pinned carabid specimens from the NEON Pu’u Maka’ala ecological observatory site in Hawaii. The collection was assembled from 600 original source images: 420 bulk images captured with a Canon EOS DSLR camera (model 7D with a 24-105 macro lens) and 180 high-detail microscopic images acquired using a SWIFCAM SC1603 system featuring a 16MP 1/2.33” CMOS sensor. The bulk images each contained 3-5 pinned specimens arranged vertically in sequential order according to their unique specimen IDs, with all individuals within a given image representing the same species and captured from the same pitfall trap. To ensure the images are optimized for advanced ecological applications like automated trait extraction, we follow the guidelines outlined in [13], which emphasize standardized specimen positioning, consistent size and color calibration, and comprehensive metadata documentation. Then we run Grounding DINO [32] to precisely detect and crop out individual beetles from these group images. The group images include comprehensive metadata comprising geolocation coordinates, collection dates, and taxonomic authentication by carabid specialists. Using the coordinates and date, we extract relevant weather data for each specimen, providing ecological context for morphological analyses. Morphological traits are measured using a digital annotation tool named TORAS [27]. Figure 5 displays a group image of carabid specimens and their corresponding individual crops.
@@
This dataset, also digitizing NEON specimens, significantly expands geographic and taxonomic coverage, comprising 11,399 images collected from 30 NEON sites across the continental United States. Unlike traditional pinned specimens, this dataset focuses on beetles preserved in ethanol-filled vials—specifically, the ‘excess’ specimens (those beyond the first 10 individuals per pitfall trap, which are pinned separately). During digitization, specimens are carefully air-dried to remove residual ethanol, mounted on minute staging sticks to standardize orientation, and imaged in bulk. However, due to the delicate nature of ethanol-preserved specimens, some individuals could not be repositioned, resulting in orientation variability. As with BeetlePUUM, we run Grounding DINO to isolate individual beetles from group images; incorporated with weather data and morphological traits (for this dataset, we use another digital annotation platform, Zooniverse [50]). Figure 6 shows a group image and individual images after detection and cropping.
@@
[47] presents a compilation of datasets BeetlePUUM and BeetlePalooza, featuring meticulously-measured morphological traits of individual specimens and offering a trait-based foundation for exploring taxonomic relationships and ecological variation among carabids. Beyond that, it serves as a valuable testbed for small-data regimes with multimodal data, a relatively less-explored area in ML where conventional approaches often underperform [56, 52].
@@ A.2.1 Summary Statistics and Distributional Insights
[…]
@@ A.2.2 Quartile Distribution
Table A.2.2 provides quartile distributions for both species and genera, illustrating the spread and central tendencies. For species, NHM-Carabids stands out with a median (Q2) of 170.0 and a third quartile (Q3) of 282.50, reflecting a higher baseline of samples per species. BeetlePUUM, despite its small species count, has a median of 32.0 and a Q3 of 62.0, indicating dense sampling. In contrast, I1MC has a median of 6.0 and a Q3 of 16.0, showing that 75% of its species have 16 or fewer samples. BeetlePalooza’s median is 35.0, but its Q4 (maximum) reaches 1,568, highlighting a long tail. For genera, NHM-Carabids again shows a high median (355.0) and Q3 (697.0), while BeetlePUUM’s median is 333.5, reflecting its focused but well-sampled genera. I1MC and BeetlePalooza have medians of 25.5 and 60.5, respectively, with maximum values (2,457 and 2,242) indicating the presence of a few heavily sampled genera.
@@ B.2 Benchmarking
Our comprehensive evaluation of vision and vision language models reveals significant performance patterns. Tables 8 and 9 present the micro and macro accuracy scores respectively, for genus and species classification across all datasets, while Table 10 provides a breakdown of multiple performance metrics. From the scores, we see that vision language models consistently outperform other approaches, with ViLT demonstrating superior performance across all datasets and metrics. ViLT achieves perfect genus-level accuracy (1.0) and exceptional species identification (0.997) on smaller, curated collections like BeetlePUUM, with corresponding perfect MCC scores (1.0 for genus, 0.995 for species). It maintains decent performance even on the challenging I1MC dataset (0.891 genus, 0.763 species micro-accuracy; MCC scores of 0.889 and 0.763 respectively), confirming that the integration of visual and textual features provides powerful taxonomic discrimination capabilities. Among other vision-language models, BioCLIP consistently ranks second, showing particularly strong performance on curated datasets but experiencing a more significant performance drop on larger, more heterogeneous collections. CLIP and SigLIP follow similar patterns but with lower performance scores. Model’s performance generally declines as the dataset size and heterogeneity increase, with all models showing a marked reduction in species-level identification accuracy on larger datasets. For instance, while ViLT maintains high precision and recall (both $>$0.99) for both genus and species on BeetlePUUM and BeetlePalooza, these metrics decline to approximately 0.74 for species classification on the I1MC dataset. Among vision-only models, supervised approaches (particularly BeIT and ConvNeXt) outperform self-supervised alternatives. BeIT achieves the best results in its category (0.923 genus, 0.821 species micro-accuracy on the merged dataset; MCC scores of 0.919 and 0.821), indicating that representations pretrained on general image collections transfer effectively to specialized taxonomic tasks. Within self-supervised models, DINOv2 consistently leads (with MCC scores reaching 0.968 for genus classification on BeetlePUUM), though it falls short of both vision-language models and supervised vision models. The performance gap between genus and species classification widens considerably in larger datasets, highlighting the increasing difficulty of fine-grained classification as taxonomic specificity increases. This pattern is consistent across all model types, with species-level F1 scores typically 10-30% points lower than genus-level scores on the larger datasets. Macro-accuracy scores reveal similar patterns, but all models show considerably lower macro-accuracy compared to micro-accuracy, particularly for species-level classification, highlighting significant class imbalance challenges in long-tailed datasets. This disparity is most pronounced in larger, more diverse datasets like I1MC, where even the top-performing ViLT model shows a substantial gap between micro-accuracy (0.763) and macro-accuracy (0.546) for species classification. The gap between genus and species classification is further amplified in macro-accuracy metrics. For example, on the merged dataset, ViLT achieves a genus macro-accuracy of 0.783 but drops to 0.657 for species classification, indicating that models struggle particularly with rare or underrepresented species. This pattern holds across all categories but is most severe for vision-only self-supervised models, where DINOv2’s species macro-accuracy reaches only 0.391 on the merged dataset.
@@ Appendix C Multi-Modal Feature Integration
We evaluate four models - BioCLIP, ConvNeXt, DINOv2, and ViLT - across three data configurations: image-only, image and morphological traits (image+traits), and image, traits, and environmental data (image+traits+env). Summary of model performance on two data sets with four models across all modality configurations is presented in Table 13. Scores show that for the 1,000-specimen subset, vision-only models showed mixed responses to additional modalities. DINOv2’s accuracy was 0.7750 with just images, dropping to 0.7600 with traits and further to 0.7550 with traits and environmental data, suggesting extra modalities were not helpful. ConvNeXt started at 0.8150 with images alone, improved slightly to 0.8350 with traits, but fell to 0.8000 with environmental data added, indicating inconsistent benefits. In contrast, vision-language models behaved differently. BioCLIP’s accuracy rose steadily from 0.8150 (image-only) to 0.8300 (image+traits) and 0.8450 (image+traits+env), showing consistent gains. ViLT, however, achieved a strong 0.9350 with images alone but remained unchanged with traits (0.9350) and dropped substantially to 0.9050 with environmental data, suggesting additional modalities may disrupt its performance. On the full dataset, trends shifted. Vision-only models benefited more from multi-modal inputs at scale. DINOv2’s accuracy increased from 0.9496 (image-only) to 0.9478 (image+traits) and 0.9513 (image+traits+env), while ConvNeXt improved from 0.9531 (image-only) to 0.9566 (image+traits) and 0.9649 (image+traits+env), indicating that additional modalities became helpful with more data. For vision-language models, BioCLIP again showed steady improvement, rising from 0.9373 (image-only) to 0.9417 (image+traits) and 0.9579 (image+traits+env), reinforcing its ability to leverage extra data. ViLT, starting near-perfect at 0.9982 (image-only), dropped marginally to 0.9956 with both traits and traits+env, suggesting limited or negative impact from additional modalities. These results reveal distinct patterns. DINOv2 and ConvNeXt struggle to benefit from extra modalities in the subset but improve at full scale, possibly due to better generalization with larger data. BioCLIP consistently gains from multi-modal inputs across both scales, highlighting its robustness. ViLT, however, shows no benefit in the subset where it suffers a substantial drop, and a marginal decline at scale, possibly indicating saturation or sensitivity to non-visual data. Given these inconsistencies, we cannot draw a firm conclusion on the effectiveness of multi-modal integration. Further experiments, varying dataset sizes, modalities, and model architectures, are needed to clarify these trends and determine optimal strategies for taxonomic classification.
@@ Appendix D Feature Mapping
[…]
@@
[…]

### v3

@@ 2 Related Work
The computational identification of insects has emerged as a critical research domain, with numerous studies addressing the challenges of taxonomic classification through advanced machine learning (ML) techniques [41, 61, 3, 63, 59, 6, 29]. Domain adaptation is critical [65, 53, 30, 18, 39]. For example, [53] investigated the generalization of ML models across different imaging conditions, revealing significant performance drops for out of distribution data. Several works proposed multimodal approaches integrating visual features with ecological metadata [13, 25, 2, 49], showing different signal modalities often evolve independently to serve distinct functions. A few recent works have explored automated methods for identifying and monitoring carabids [7, 26, 23, 33, 50, 62, 1]. Most notable among them is NEON’s continental-scale ground beetle monitoring program [26, 5, 42, 44, 43]. Among others, [7] achieved 84.6% species-level accuracy using linear discriminant analysis on pitfall trap-collected carabids; [23] achieved 51.9% species-level accuracy using applied neural networks; and [1] got 90% accuracy on Sri Lankan beetles by using background removal, segmentation, and transfer learning. However, these approaches are limited by small taxonomic coverage and the use of conventional ML models that lack the representational power of cutting-edge vision encoders. Our work addresses these limitations by leveraging that power and covering a broader taxonomic range.
@@ Acknowledgment
This work was supported by both the Imageomics Institute and the AI and Biodiversity Change (ABC) Global Center. The Imageomics Institute is funded by the US National Science Foundation’s Harnessing the Data Revolution (HDR) program under Award No. 2118240 (Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning). The ABC Global Center is funded by the US National Science Foundation under Award No. 2330423 and Natural Sciences and Engineering Research Council of Canada under Award No. 585136. This work draws on research funded by the Social Sciences and Humanities Research Council. S. R. and A. E. were additionally supported by the US National Science Foundation’s Award No. 242918 (EPSCOR Research Fellows: NSF: Advancing National Ecological Observatory Network-Enabled Science and Workforce Development at the University of Maine with Artificial Intelligence) and by Hatch project Award #MEO-022425 from the US Department of Agriculture’s National Institute of Food and Agriculture. This work draws on research supported by the National Ecological Observatory Network (NEON), a program sponsored by the National Science Foundation and operated under cooperative agreement by Battelle. This material uses specimens and/or samples collected as part of the NEON Program and is based in part upon work supported by NEON. Additionally, this material is based upon work supported by the National Science Foundation under Award Numbers 2301322, 1950364, and 1926569. This publication uses data generated via the Zooniverse platform, the development of which is funded by generous support, including a Global Impact Award from Google, and by a grant from the Alfred P. Sloan Foundation.
@@
All four datasets used in this study are publicly available: BeetlePUUM [^1] Official name on Hugging Face: imageomics/Hawaii-beetles [51], BeetlePalooza [^2] Official name on Hugging Face: imageomics/2018-NEON-beetles [17], NHM-Carabids [22], and I1MC [^3] Official name: uark-cviu.github.io/projects/insect-foundation [45].
@@
All code developed for and utilized in this study is publicly available on github at Imageomics/BeetleVerse.
Competing Interests.
The authors declare no competing interests.
@@ A.1 Dataset Details
This dataset digitizes all pinned carabid specimens from the NEON Pu’u Maka’ala ecological observatory site in Hawaii [51]. The specimens originate from the NEON biorepository carabid pinned voucher collection curated by the Bernice Pauahi Bishop Museum [5], with associated specimen- and trap-level metadata provided through the NEON ground beetle pitfall sampling data product [42]. The dataset is publicly available on HuggingFace, at imageomics/Hawaii-beetles. The collection was assembled from 600 original source images: 420 bulk images captured with a Canon EOS DSLR camera (model 7D with a 24–105 macro lens) and 180 high-detail microscopic images acquired using a SWIFCAM SC1603 system featuring a 16MP 1/2.33” CMOS sensor. The bulk images each contained 3–5 pinned specimens arranged vertically in sequential order according to their unique specimen IDs, with all individuals within an image representing the same species and captured from the same pitfall trap.
To ensure the images are optimized for advanced ecological applications such as automated trait extraction, the curators of the dataset follow the digitization guidelines outlined in [14], which emphasize standardized specimen positioning, consistent size and color calibration, and comprehensive metadata documentation. Grounding DINO [34] is then applied to precisely detect and crop individual beetles from the group images. The original group images include comprehensive metadata comprising geolocation coordinates, collection dates, and taxonomic authentication by carabid specialists. Using the collection coordinates and dates, relevant weather data are extracted for each specimen to provide ecological context for morphological analyses. Morphological traits are measured using the TORAS digital annotation tool [28]. Figure 5 displays a representative group image and the corresponding individual specimen crops.
@@
This dataset digitizes NEON carabid specimens and substantially expands both geographic and taxonomic coverage, comprising 11,399 images collected from 30 NEON sites across the continental United States [17]. The specimens are drawn from the NEON biorepository ethanol-preserved collections, associated with trap sorting and archive pooling workflows [44, 43]. Unlike traditional pinned voucher specimens, this dataset focuses on excess individuals preserved in ethanol-filled vials — specifically, specimens collected beyond ten individuals per pitfall trap, which are pinned separately under NEON protocols. The dataset is publicly available on HuggingFace, at imageomics/2018-NEON-beetles.
During digitization, specimens are carefully air-dried to remove residual ethanol, mounted on minute staging sticks to standardize orientation, and imaged in bulk. Due to the delicate nature of ethanol-preserved specimens, some individuals could not be repositioned, resulting in variability in specimen orientation. As with BeetlePUUM, Grounding DINO [34] is applied to isolate individual beetles from group images. Each specimen image is linked with NEON-provided collection metadata and supplemented with site-level weather data derived from collection date and geolocation. Morphological traits for this dataset are annotated using the Zooniverse digital annotation platform [54]. Figure 6 presents a representative group image and the corresponding individual specimen crops produced through the detection and cropping pipeline.
@@
[50] presents a compilation of these two datasets, featuring meticulously-measured morphological traits of individual specimens and offering a trait-based foundation for exploring taxonomic relationships and ecological variation among carabids. Moreover, it serves as a valuable testbed for small-data regimes with multimodal data, a relatively less-explored area in ML where conventional approaches often underperform [60, 55].
@@ A.2.1 Summary Statistics and Distributional Insights
Dataset Mean Median Std Dev Min Q1 (25%) Q3 (75%) IQR Max Total Species Total Samples Skewness Kurtosis
@@ A.2.2 Quartile Distribution
Table 5 provides quartile distributions for both species and genera, illustrating the spread and central tendencies. For species, NHM-Carabids stands out with a median (Q2) of 170.0 and a third quartile (Q3) of 282.50, reflecting a higher baseline of samples per species. BeetlePUUM, despite its small species count, has a median of 32.0 and a Q3 of 62.0, indicating dense sampling. In contrast, I1MC has a median of 6.0 and a Q3 of 16.0, showing that 75% of its species have 16 or fewer samples. BeetlePalooza’s median is 35.0, but its Q4 (maximum) reaches 1,568, highlighting a long tail. For genera, NHM-Carabids again shows a high median (355.0) and Q3 (697.0), while BeetlePUUM’s median is 333.5, reflecting its focused but well-sampled genera. I1MC and BeetlePalooza have medians of 25.5 and 60.5, respectively, with maximum values (2,457 and 2,242) indicating the presence of a few heavily sampled genera.
@@ B.2 Benchmarking
Our comprehensive evaluation of vision and vision language models reveals significant performance patterns. Tables 8 and 9 present the micro and macro accuracy scores respectively, for genus and species classification across all datasets, while Table 10 provides a breakdown of multiple performance metrics. From the scores, we see that vision language models consistently outperform other approaches, with ViLT demonstrating superior performance across all datasets and metrics. ViLT achieves perfect genus-level accuracy (1.0) and exceptional species identification (0.997) on smaller, curated collections like BeetlePUUM, with corresponding perfect MCC scores (1.0 for genus, 0.995 for species). It maintains decent performance even on the challenging I1MC dataset (0.891 genus, 0.763 species micro-accuracy; MCC scores of 0.889 and 0.763 respectively), confirming that the integration of visual and textual features provides powerful taxonomic discrimination capabilities. Among other vision-language models, BioCLIP consistently ranks second, showing particularly strong performance on curated datasets but experiencing a more significant performance drop on larger, more heterogeneous collections. CLIP and SigLIP follow similar patterns but with lower performance scores.
Model’s performance generally declines as the dataset size and heterogeneity increase, with all models showing a marked reduction in species-level identification accuracy on larger datasets. For instance, while ViLT maintains high precision and recall (both >>0.99) for both genus and species on BeetlePUUM and BeetlePalooza, these metrics decline to approximately 0.74 for species classification on the I1MC dataset. Among vision-only models, supervised approaches (particularly BeIT and ConvNeXt) outperform self-supervised alternatives. BeIT achieves the best results in its category (0.923 genus, 0.821 species micro-accuracy on the merged dataset; MCC scores of 0.919 and 0.821), indicating that representations pretrained on general image collections transfer effectively to specialized taxonomic tasks. Within self-supervised models, DINOv2 consistently leads (with MCC scores reaching 0.968 for genus classification on BeetlePUUM), though it falls short of both vision-language models and supervised vision models. The performance gap between genus and species classification widens considerably in larger datasets, highlighting the increasing difficulty of fine-grained classification as taxonomic specificity increases. This pattern is consistent across all model types, with species-level F1 scores typically 10-30% points lower than genus-level scores on the larger datasets.
Macro-accuracy scores reveal similar patterns, but all models show considerably lower macro-accuracy compared to micro-accuracy, particularly for species-level classification, highlighting significant class imbalance challenges in long-tailed datasets. This disparity is most pronounced in larger, more diverse datasets like I1MC, where even the top-performing ViLT model shows a substantial gap between micro-accuracy (0.763) and macro-accuracy (0.546) for species classification. The gap between genus and species classification is further amplified in macro-accuracy metrics. For example, on the merged dataset, ViLT achieves a genus macro-accuracy of 0.783 but drops to 0.657 for species classification, indicating that models struggle particularly with rare or underrepresented species. This pattern holds across all categories but is most severe for vision-only self-supervised models, where DINOv2’s species macro-accuracy reaches only 0.391 on the merged dataset.
@@ Appendix C Multi-Modal Feature Integration
We evaluate four models - BioCLIP, ConvNeXt, DINOv2, and ViLT - across three data configurations: image-only, image and morphological traits (image+traits), and image, traits, and environmental data (image+traits+env). Summary of model performance on two data sets with four models across all modality configurations is presented in Table 13. Scores show that for the 1,000-specimen subset, vision-only models showed mixed responses to additional modalities. DINOv2’s accuracy was 0.7750 with just images, dropping to 0.7600 with traits and further to 0.7550 with traits and environmental data, suggesting extra modalities were not helpful. ConvNeXt started at 0.8150 with images alone, improved slightly to 0.8350 with traits, but fell to 0.8000 with environmental data added, indicating inconsistent benefits. In contrast, vision-language models behaved differently. BioCLIP’s accuracy rose steadily from 0.8150 (image-only) to 0.8300 (image+traits) and 0.8450 (image+traits+env), showing consistent gains. ViLT, however, achieved a strong 0.9350 with images alone but remained unchanged with traits (0.9350) and dropped substantially to 0.9050 with environmental data, suggesting additional modalities may disrupt its performance.
On the full dataset, trends shifted. Vision-only models benefited more from multi-modal inputs at scale. DINOv2’s accuracy increased from 0.9496 (image-only) to 0.9478 (image+traits) and 0.9513 (image+traits+env), while ConvNeXt improved from 0.9531 (image-only) to 0.9566 (image+traits) and 0.9649 (image+traits+env), indicating that additional modalities became helpful with more data. For vision-language models, BioCLIP again showed steady improvement, rising from 0.9373 (image-only) to 0.9417 (image+traits) and 0.9579 (image+traits+env), reinforcing its ability to leverage extra data. ViLT, starting near-perfect at 0.9982 (image-only), dropped marginally to 0.9956 with both traits and traits+env, suggesting limited or negative impact from additional modalities. These results reveal distinct patterns. DINOv2 and ConvNeXt struggle to benefit from extra modalities in the subset but improve at full scale, possibly due to better generalization with larger data. BioCLIP consistently gains from multi-modal inputs across both scales, highlighting its robustness. ViLT, however, shows no benefit in the subset where it suffers a substantial drop, and a marginal decline at scale, possibly indicating saturation or sensitivity to non-visual data. Given these inconsistencies, we cannot draw a firm conclusion on the effectiveness of multi-modal integration. Further experiments, varying dataset sizes, modalities, and model architectures, are needed to clarify these trends and determine optimal strategies for taxonomic classification.
@@ Appendix D Feature Mapping
Case Train Test Type #Taxa Model Accuracy Precision Recall F1-Score MCC NHM-I1M-genus NHM I1M genus 57 BioCLIP 0.3899 0.6190 0.3899 0.4103 0.3623 NHM-I1M-genus NHM I1M genus 57 CLIP 0.1946 0.3595 0.1946 0.1844 0.1425 NHM-I1M-genus NHM I1M genus 57 ConvNeXt 0.3603 0.5438 0.3603 0.3610 0.3205 NHM-I1M-genus NHM I1M genus 57 LeViT 0.2497 0.3919 0.2497 0.2510 0.2070 NHM-I1M-genus NHM I1M genus 57 SWINv2 0.3238 0.4822 0.3238 0.3165 0.2898 NHM-I1M-genus NHM I1M genus 57 ViLT 0.6907 0.8168 0.6907 0.6966 0.6736 NHM-I1M-species NHM I1M species 68 BioCLIP 0.4221 0.6546 0.4221 0.4362 0.4133 NHM-I1M-species NHM I1M species 68 CLIP 0.0875 0.2436 0.0875 0.0856 0.0761 NHM-I1M-species NHM I1M species 68 ConvNeXt 0.2291 0.4589 0.2291 0.2507 0.2204 NHM-I1M-species NHM I1M species 68 LeViT 0.1120 0.1942 0.1120 0.1195 0.1026 NHM-I1M-species NHM I1M species 68 SWINv2 0.1750 0.3112 0.1750 0.1624 0.1618 NHM-I1M-species NHM I1M species 68 ViLT 0.5740 0.7737 0.5740 0.6132 0.5680 BPZ-I1M-genus BPZ I1M genus 33 BioCLIP 0.3553 0.5198 0.3553 0.3558 0.3257 BPZ-I1M-genus BPZ I1M genus 33 CLIP 0.1386 0.2623 0.1386 0.1489 0.1071 BPZ-I1M-genus BPZ I1M genus 33 ConvNeXt 0.3464 0.4843 0.3464 0.3430 0.3142 BPZ-I1M-genus BPZ I1M genus 33 LeViT 0.1985 0.3115 0.1985 0.2042 0.1725 BPZ-I1M-genus BPZ I1M genus 33 SWINv2 0.3395 0.3881 0.3395 0.3202 0.2986 BPZ-I1M-genus BPZ I1M genus 33 ViLT 0.6001 0.6931 0.6001 0.5823 0.5756 BPZ-I1M-species BPZ I1M species 72 BioCLIP 0.3656 0.4298 0.3656 0.3400 0.3558 BPZ-I1M-species BPZ I1M species 72 CLIP 0.1128 0.2835 0.1128 0.1107 0.1025 BPZ-I1M-species BPZ I1M species 72 ConvNeXt 0.2592 0.3895 0.2592 0.2394 0.2498 BPZ-I1M-species BPZ I1M species 72 LeViT 0.1422 0.1897 0.1422 0.1413 0.1276 BPZ-I1M-species BPZ I1M species 72 SWINv2 0.1913 0.3177 0.1913 0.1855 0.1802 BPZ-I1M-species BPZ I1M species 72 ViLT 0.4757 0.4998 0.4757 0.4287 0.4676 NHM-BPZ-genus NHM BPZ genus 16 BioCLIP 0.4632 0.7094 0.4632 0.5178 0.4222 NHM-BPZ-genus NHM BPZ genus 16 CLIP 0.3076 0.4796 0.3076 0.2993 0.2161 NHM-BPZ-genus NHM BPZ genus 16 ConvNeXt 0.3697 0.6106 0.3697 0.3770 0.3010 NHM-BPZ-genus NHM BPZ genus 16 LeViT 0.3481 0.4962 0.3481 0.3284 0.2591 NHM-BPZ-genus NHM BPZ genus 16 SWINv2 0.4371 0.5217 0.4371 0.3886 0.3546 NHM-BPZ-genus NHM BPZ genus 16 ViLT 0.9230 0.9552 0.9230 0.9311 0.9106
@@
Dataset Images Data Type Model Acc Prec Rec F1 MCC Subset 1000 image BioCLIP 0.8150 0.7585 0.8150 0.7780 0.8038 Subset 1000 image ConvNeXt 0.8150 0.7733 0.8150 0.7864 0.8037 Subset 1000 image DINOv2 0.7750 0.7313 0.7750 0.7349 0.7609 Subset 1000 image ViLT 0.9350 0.9121 0.9350 0.9172 0.9314 Subset 1000 image+traits BioCLIP 0.8300 0.7603 0.8300 0.7850 0.8198 Subset 1000 image+traits ConvNeXt 0.8350 0.7899 0.8350 0.8052 0.8251 Subset 1000 image+traits DINOv2 0.7600 0.7312 0.7600 0.7290 0.7449 Subset 1000 image+traits ViLT 0.9350 0.9121 0.9350 0.9172 0.9314 Subset 1000 image+traits+env BioCLIP 0.8450 0.7958 0.8450 0.8064 0.8347 Subset 1000 image+traits+env ConvNeXt 0.8000 0.7304 0.8000 0.7528 0.7863 Subset 1000 image+traits+env DINOv2 0.7550 0.6896 0.7550 0.7030 0.7379 Subset 1000 image+traits+env ViLT 0.9050 0.8783 0.9050 0.8814 0.8990 Full 11372 image BioCLIP 0.9373 0.9325 0.9373 0.9323 0.9330 Full 11372 image ConvNeXt 0.9531 0.9426 0.9531 0.9461 0.9498 Full 11372 image DINOv2 0.9496 0.9453 0.9496 0.9446 0.9461 Full 11372 image ViLT 0.9982 0.9974 0.9982 0.9978 0.9981 Full 11372 image+traits BioCLIP 0.9417 0.9357 0.9417 0.9368 0.9375 Full 11372 image+traits ConvNeXt 0.9566 0.9512 0.9566 0.9514 0.9535 Full 11372 image+traits DINOv2 0.9478 0.9445 0.9478 0.9430 0.9441 Full 11372 image+traits ViLT 0.9956 0.9951 0.9956 0.9948 0.9953 Full 11372 image+traits+env BioCLIP 0.9579 0.9531 0.9579 0.9536 0.9549 Full 11372 image+traits+env ConvNeXt 0.9649 0.9604 0.9649 0.9604 0.9624 Full 11372 image+traits+env DINOv2 0.9513 0.9502 0.9513 0.9468 0.9479 Full 11372 image+traits+env ViLT 0.9956 0.9952 0.9956 0.9950 0.9953